OpenStudy (anonymous):

For y = x^2 + 6x − 16, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation. Show all work and use complete sentences to receive full credit.

OpenStudy (anonymous):

i think its an up and minimum but that's all i know

OpenStudy (anonymous):

anytime a is greater than 0 in a quadratic the parabola is going to open 'up'. If a is negative it will be reflected across the x axis and open 'down'

OpenStudy (anonymous):

tyvm

OpenStudy (anonymous):

and yes it is a minimum

OpenStudy (anonymous):

the rest of it should be on the link nurali provided

OpenStudy (anonymous):

but just incase y = x^2 + 6x − 16 y + 9 = (x^2 + 6x + 9) -16 y+9=(x+3)^2-16 y=(x+3)^2- 25 so the vertex is going to be -3,-25

OpenStudy (anonymous):

and what i just did is called completing the square. you take the b value(6) divide it by 2, then square it and add it to both sides.

OpenStudy (anonymous):

then group and factor the new quadratic like normal, and subtract the new number from both sides to isolate y

OpenStudy (anonymous):

the x intercepts will 8 and -2 since y = x^2 + 6x − 16 = y = (x+8)(x-2)

OpenStudy (anonymous):

to find those you just factor the original quadratic like usual. i don't really understand what is meant by describe the graph but i think they mean that the graph is translated 3 units to the left and 25 units down